OpenStudy (anonymous):

are either of these polynomials? *f(x)=x^3-11x^2+3^x *f(x)=2x^3-5x^5-2/9x^2+9

OpenStudy (anonymous):

not sure if a polynomial can a variable for an exponent

OpenStudy (anonymous):

A polynomial is a polynomial if and only if the co-efficients are all real and all the exponents of the terms are real numbers and positive. So for example, \(f(x)=x^3-11x^2+3^x\) Is NOT a polynomial because the third term has a exponent which is not a real number, But a variable.

OpenStudy (anonymous):

Just a question, is the second one, \(f(x)=2x^3-5x^5-\frac{2}{9x^2}+9\) ?

OpenStudy (anonymous):

what about if there was a fraction with a variable on the side?

OpenStudy (anonymous):

like the "co-efficient" is the fraction, or the "exponent" is the fraction or the variable is the fraction?

OpenStudy (anonymous):

like for ex: (2/9)x^2 .

OpenStudy (anonymous):

That is OKAY. Because \(\frac{2}{9}\), Though it is a fraction represents a number (0.22222...) So it is a valid co-efficient.

OpenStudy (anonymous):

Think about it like this. Would \((\frac{2}{9}x^2)\) Look more "polynomial-ish" if it was \((0.222x^2)\)?

OpenStudy (anonymous):

cool thank you!